-y - 4x = 14
y = -5x - 15

Answer by Cromlix(4381)

Hi there,
-y - 4x = 14
y = -5x - 15
.....
-y - 4x = 14
y + 5x = -15
Add
........
x = -1
Substitute x = -1 in:
y = -5x -15
y = -5(-1) - 15
y = 5 - 15
y = -10
{-1,-10}
Hope this helps:-)